OpenStudy (anonymous):

what are the points where they intersect x-2y=-2 and 3x-2y=4

OpenStudy (anonymous):

you gotta right the equations as y= ax+b and y_{1}=\[a _{1}x+b_{1}\] then sove y=y_{1} for your example: x-2y=4 and 3x-2y_{1}=4 <> x-4=2y 3x-4=2y_{1} <> y=x/2 -2 3x/2 -2=y_{1} solve y=y_{1}

OpenStudy (anonymous):

is it 4,4

OpenStudy (anonymous):

x-2y=-4 not -2

OpenStudy (precal):

(3, 2.5) is the point of intersection

OpenStudy (anonymous):

and its with the x-2y=-4

OpenStudy (anonymous):

then x=0